# Universality conditions of unified classical and quantum reservoir computing

### v2

@@ 4 Quantum echo state networks and universality
The last Section applies the concept of universality to quantum reservoirs based on qubit registers (Fujii and Nakajima (2017); Nakajima et al. (2019)). Such extension exploits the density matrix representation of a quantum state and involves spatial multiplexing. An effective implementation of provably universal quantum reservoir systems, exploiting dissipative quantum systems, was proposed in Refs. Chen and Nurdin (2019); Chen et al. (2020).
@@ 4.1 Quantum echo state networks
Let $\mathcal{H}$ be a Hilbert space, that we assume to be finite-dimensional for the sake of simplicity, and let $\mathcal{B}(\mathcal{H})$ be the space of bounded operators on $\mathcal{H}$, equipped with the Schatten norm $\norm{T}=\text{Tr}(A^{\dagger})$. Any pure or mixed state in $\mathcal{H}$ may be represented by a density operator $\rho\in\mathcal{B}(\mathcal{H})$, namely, by an element in the compact convex subspace
@@
At each time step, the reservoir state is represented by a density operator defined on a suitable Hilbert space. The quantum dynamics of the reservoir is described by a quantum channel (see Weedbrook et al. (2012) for a detailed introduction of this concept).
@@
for some $r<1$, as a consequence of Lemma 3.3. In this regard, we recall that any CPTP map is non-expansive, meaning that Eq. 10 holds with $r\leq 1$ (see Theorem 9.2 in Nielsen and Chuang (2010)). With the same argument developed in Lemma 3.2, one concludes that a qESN with echo state property falls in our definition of echo reservoir computer.
@@ 4.2 A class of quantum reservoir computers
In these last two subsections, we present an embodiment of a qESN exploiting a $N$-qubit register proposed in Nakajima et al. (2019). Let $\mathcal{H}_{2}$ be the two dimensional complex Hilbert space associated with a single qubit and $\mathcal{H}^{N}=\bigotimes_{j=1}^{N}\mathcal{H}_{2}$ the $2^{N}$-dimensional Hilbert space of the whole register. Denoting with $I=\sigma_{00},\,X=\sigma_{01},\,Y=\sigma_{10},Z=\sigma_{11}$ the Pauli operators, the products $\left\{P_{i}\right\}=\left\{\bigotimes_{l=1}^{N}\sigma_{i_{2k-1}i_{2k}}\right\}$ form a basis of the normed space $\mathcal{B}(\mathcal{H}^{N})$. Then, the density operator $\rho$ that describes the state of the reservoir is represented as a $4^{N}$-dimensional vector $r=\left(r^{00\dots 00},r^{01\dots 00},\dots,r^{11\dots 11}\right)$ with
@@
ESN Time series in $\mathbb{R}^{n}$ Artificial neural network $T=\mathbb{Z}$ Linear or polynomial function of every node in the reservoir Jaeger (2001); Grigoryeva and Ortega (2018b, a),
LSM Spike trains Spiking neural network $T=\mathbb{R}$ Analogical function of a subset of neurons Maass et al. (2002)
qESN Time series in $\mathbb{R}^{n}$ Qubit register $T=\mathbb{Z}$ Linear function of the $Z$ Pauli operators Nakajima et al. (2019); Chen et al. (2020)
ERC Orbits in any normed space Any system described by a normed space $T=\mathbb{Z}$ or $T=\mathbb{R}$ Generic functional applied to the reservoir state This work
@@ 4.3 Universality by spatial multiplexing
Following the result in Theorem 2.15, we show how spatial multiplexity ensures the universality of a class of qESN generated from the systems described in Subsection 4.2, as it allows to represent sums and products of functionals. Spatial multiplexing consists of preparing uncoupled instances of a reservoir system with different internal parameters and running them in parallel with the same input. More formally, let $\mathcal{Q}_{(H_{1},w)},\mathcal{Q}_{(H_{2},v)}$ be two independent qESN systems with respectively $N_{1}$ and $N_{2}$ qubit registers. Denote with $R_{1}$ and $R_{2}$ the respective associated functionals, and with $r^{1}$ and $r^{2}$ the respective reservoir states represented in the Pauli basis. Spatial multiplexity provides the construction of the system whose associated functional is the product of $R_{1}$ and $R_{2}$. The reservoir state $r$ of the spatial multiplexed system is the joint reservoir state of each isolated register, namely $r=(r^{1},r^{2})$. Consequently, we denote the true nodes, corresponding to the $Z$ operators of the two registers, respectively as $z^{1},\dots,z^{N_{1}}$ and $z^{N_{1}+1},\dots,z^{N_{1}+N_{2}}$. Then, the functional associated with the qESN
@@
The sum is built identically by exploiting $\sum_{i=1}^{N_{1}}w_{i}z_{t+1}^{i}+\sum_{i={N_{1}+1}}^{N_{2}}v_{i}z_{t+1}^{i}$ as readout. As long as the unitary dynamics induces fading memory and is able to separate different inputs, we can conclude that the family of quantum reservoir computers that are spatial multiplexed instances of systems described in Eq.(11) is universal.

### v3

@@ 4 Quantum echo state networks and universality
The last Section applies the concept of universality to quantum reservoirs based on qubit registers [48, 1]. Such extension exploits the density matrix representation of a quantum state and involves spatial multiplexing. An effective implementation of provably universal quantum reservoir systems, exploiting dissipative quantum systems, was proposed in Refs. [60, 61].
@@ 4.1 Quantum echo state networks
Let $\mathcal{H}$ be a Hilbert space, that we assume to be finite-dimensional for the sake of simplicity, and let $\mathcal{B}(\mathcal{H})$ be the space of bounded operators on $\mathcal{H}$, equipped with the Schatten norm $\norm{T}=\text{Tr}(A^{\dagger}A)$. Any pure or mixed state in $\mathcal{H}$ may be represented by a density operator $\rho\in\mathcal{B}(\mathcal{H})$, namely, by an element in the compact convex subspace
@@
At each time step, the reservoir state is represented by a density operator defined on a suitable Hilbert space. The quantum dynamics of the reservoir is described by a quantum channel (see [68] for a detailed introduction of this concept).
@@
for some $r<1$, as a consequence of Lemma 3.3. In this regard, we recall that any CPTP map is non-expansive, meaning that Eq. 10 holds with $r\leq 1$ (see Theorem 9.2 in [69]). With the same argument developed in Lemma 3.2, one concludes that a qESN with echo state property falls in our definition of echo reservoir computer.
@@ 4.2 A class of quantum reservoir computers
In these last two subsections, we briefly present an embodiment of a qESN exploiting a $N$-qubit register proposed in [1] and revised in [60, 61]. Let $\mathcal{H}_{2}$ be the two-dimensional complex Hilbert space associated with a single qubit and $\mathcal{H}^{N}=\bigotimes_{j=1}^{N}\mathcal{H}_{2}$ the $2^{N}$-dimensional Hilbert space of the whole register. Denoting with $I=\sigma_{00},\,X=\sigma_{01},\,Y=\sigma_{10},Z=\sigma_{11}$ the Pauli operators, the products $\left\{P_{i}\right\}=\left\{\bigotimes_{l=1}^{N}\sigma_{i_{2k-1}i_{2k}}\right\}$ form a basis of the normed space $\mathcal{B}(\mathcal{H}^{N})$. Then, the density operator $\rho$ that describes the state of the reservoir is represented by a $4^{N}$-dimensional vector $r=\left(r^{00\dots 00},r^{01\dots 00},\dots,r^{11\dots 11}\right)$ with
@@
ESN Time series in $\mathbb{R}^{n}$ Artificial neural network $T=\mathbb{Z}$ Linear or polynomial function of every node in the reservoir [3, 26, 27],
LSM Spike trains Spiking neural network $T=\mathbb{R}$ Analogical function of a subset of neurons [5]
qESN Time series in $\mathbb{R}^{n}$ Qubit register $T=\mathbb{Z}$ Linear function of the $Z$ Pauli operators [1, 61]
ERC Orbits in any normed space Any system in a normed space $T=\mathbb{Z}$ or $T=\mathbb{R}$ Generic functional applied to the reservoir state This work
@@ 4.3 Universality by spatial multiplexing
Following Theorem 2.15, we show how the class of spatial multiplexed realizations of qESN generated from the systems described in Subsection 4.2 is universal, as spatial multiplexity allows the representation of sums and products of the associated functionals. Informally, spatial multiplexing consists of preparing uncoupled instances of a reservoir system with different internal parameters and running them in parallel with the same input. More formally, let $\mathcal{Q}_{(H_{1},w)},\mathcal{Q}_{(H_{2},v)}$ be two independent qESN systems with respectively $N_{1}$ and $N_{2}$ qubit registers. Denote with $R_{1}$ and $R_{2}$ the respective associated functionals, and with $r^{1}$ and $r^{2}$ the respective reservoir states represented in the Pauli basis. Spatial multiplexity provides the construction of the system whose associated functional is the product (or, similarly the sum) of $R_{1}$ and $R_{2}$. The reservoir state $r$ of the spatial multiplexed system is the joint reservoir state of each isolated register, namely $r=(r^{1},r^{2})$. Consequently, we denote the true nodes, corresponding to the $Z$ operators of the two registers, respectively as $z^{1},\dots,z^{N_{1}}$ and $z^{N_{1}+1},\dots,z^{N_{1}+N_{2}}$. Then, the functional associated with the qESN
@@
The sum is built identically by exploiting $\sum_{i=1}^{N_{1}}w_{i}z_{t+1}^{i}+\sum_{i={N_{1}+1}}^{N_{2}}v_{i}z_{t+1}^{i}$ as readout. As long as the unitary dynamics induces fading memory and is able to separate different inputs, we can conclude that the family of quantum reservoir computers that are spatial multiplexed instances of systems described in Eq.(11) is universal. In many cases, the proof of separability is a challenging issue. We refer to [61] for a detailed discussion on how separability is achieved formally.